What is the vertex of y=2x2+4x -3
Answer by Theo(13342)

the equation is y = 2x^2 + 4x - 3

this equation is in standard form, where:

a = coefficient of x^2 term.
b = coefficient of x term.
c = constant term.

the vertex is at (x = -b/2a, y = f(-b/2a))

a = 2
b = 4
c = -3

to find the value of x, use the formula:

x = -b/2a

in this equation, b = 4 and a = 2

x = -b/2a becomes x = -b/2a = -4/4 which becomes:

x = -b/2a = -1

to find f(-b/2a), replace x in the quadratic equation with -b/2a.

since -b/2a = -1, replace x with -1 in the equation to get:

y = 2x^2 + 4x - 3 becomes y = 2*(-1)^2 + 4*(-1) - 3

simplify to get y = 2 - 4 - 3

combine like terms to get y = -5

you have x = -1 and y = -5.

the vertex is at (x,y) = (-1,-5).

the following graph of the equation confirms this is true.